Define the linear function f:R3→R2 by f(x,y,z)=(x-z,y-x,z-y).Find

i) the kernel of f

ii)the nulllity of f

iii)the rank of f and a basis for the range of f


1
Expert's answer
2021-08-17T08:34:19-0400

Consider the linear function "f:\R^3→\R^3,\ f(x,y,z)=(x-z,y-x,z-y)." .


i) Let us find the kernel of "f:"

"ker f=\{(x,y,z)\in R^3,\ f(x,y,z)=(0,0,0)\}\\
=\{(x,y,z)\in R^3,\ (x-z,y-x,z-y)=(0,0,0)\}\\
=\{(x,y,z)\in R^3,\ x-z=0,y-x=0,z-y=0\}\\
=\{(x,y,z)\in R^3,\ x=y=z\}\\
=\{(x,x,x),\ x\in \R\}\\
=\{x(1,1,1),\ x\in \R\}"


ii) It follows that the vector "(1,1,1)" form a basis of the kernel, and hence the nulllity of "f" is 1.


iii) Let the vector "(a,b,c)" belongs to the range of "f." Then "f(x,y,z)=(x-z,y-x,z-y)=(a,b,c)" for some "(x,y,z)\in\R^3." It follows that "x-z=a,y-x=b,z-y=c." Then "(x-z)+(y-x)=a+b,z-y=c," that is "-z+y=a+b,z-y=c." We conclude that "a+b=-c." Since "(a,b,c)=(a,b,-(a+b))=a(1,0,-1)+b(0,1,-1)," and by the rank-nullity theorem "rank f=3-1=2," we conclude that "\{(1,0,-1),(0,1,-1)\}" is a basis of for the range "f".